# Sparse identification of nonlinear dynamics and Koopman operators with Shallow Recurrent Decoder Networks

### v1

@@ 3.2.2 Latent space regularization via ensemble SINDy
We first note that we deviate from the original SHRED architecture by using a GRU as opposed to an LSTM. This choice was made because we generally found that GRU provides a smoother latent space. Now, recall that our goal is to find a SHRED model with a latent state that is within the SINDy-class functional. However, the initial latent representation from SHRED does not follow the SINDy-based ODE structure at all. On the one hand, if we naively apply SINDy to the initial latent representation, the discovery is unlikely to fit the latent representation trajectory. On the other hand, if we directly replace the GRU unit to SINDy and force the latent space to follow the discovered SINDy model, it might lose information that is important to reconstruction the entire spatial domain. Therefore, it is important to let the two latent spaces align progressively.
In Algorithm 1, we describe our training procedure that allows the two trajectories to progressively align with each other. To further ensure a gradual adaptation and avoid over-regularization, we introduce ensemble SINDy units with varying levels of sparsity constraints, which ranges the effect from promoting a full model (all terms in the library are active) to a null model (where no dynamics are represented). From the initial latent representation ${\bm{z}}^{\text{iter 0}}_{1:t}$ from SHRED, the SINDy model first provides an initial estimate of ensemble SINDy coefficients $\{\hat{\Xi}_{0}^{i}\}_{i=b}^{B}$. Then, the parameters of SHRED will be updated towards the dynamics simulated by $\{\hat{\Xi}_{0}\}_{i=b}^{B}$, which generates a new latent representation trajectory ${\bm{z}}^{\text{iter 1}}_{1:t}$. We iterate this procedure and jointly optimize the following loss function to let the SHRED latent representation trajectory approximate the SINDy generated trajectory:
@@ 3.2.3 Latent space linearization via the Koopman operator
Koopman operator theory [27; 28] provides an alternative approach to solving these problems by linearizing the underlying dynamics. The linearized embedding is theoretically grounded to be able to represent nonlinear dynamics in a linear framework, which is desirable for many applications in science and engineering, including control, robotics, weather modeling, and so on. From the transformed measurements ${\bm{z}}=g({\bm{x}})$ from the true system ${\bm{x}}$, the Koopman operator $\mathcal{K}$ is an infinite-dimensional linear operator that
@@
To enable continuous spectrum for better approximation for the Koopman operator, one could adapt an additional neural network to learn the eigenvalues $\lambda_{i}$’s and form the linear dynamics $K$ from the learned eigenvalues [35]. We also provide detailed implementation and experimental result in the Appendix. This flexibility allows the system to learn dynamical systems in a more general setting, accommodating various initial conditions and experimental setups. In our experimental study, we observe that this strategy may perform comparably to learning a fixed ${\bm{K}}$ for linear dynamics under fixed experimental environments.
3.3 Training setup of dynamical system learning
The dynamical system we wish to model has the form that
$\dot{x}=f(x),$ (8)
@@ 3.3 Training setup of dynamical system learning
Suppose that the dynamical system has the form $\dot{{\bm{x}}}=f({\bm{x}})$, and we have measurements of ${\bm{x}}=\{{\bm{x}}_{1},{\bm{x}}_{2},...,{\bm{x}}_{t},{\bm{x}}_{t+1},...,{\bm{x}}_{T}\}$ with time gap $\Delta t$. The empirical loss function $\hat{L}(\cdot)$ of a function $f$ given the empirical distribution $P_{n}$ which consists of data samples $\mathcal{D}=\{({\bm{x}}_{i},\dot{{\bm{x}}_{i}})\}_{i=1}^{n}$ is
$\hat{L}\left(f\right)=\frac{1}{n}\sum_{i=1}^{n}\ell\left(f({\bm{x}}_{i}),{\bm{x}}_{i}\right),$ (9)
while the true loss is
$L\left(f\right)=\mathbb{E}\left[\ell\left(f(x),y\right)\right].$ (10)
@@
In practice, we only have access to data samples (or equivalently empirical distributions). There is a generalization gap between $L(f)$ and $L(\hat{f})$, which is correlated with the expressive power of the functional class. When the expressive power of the estimator functional class is small, the generalization error is expected to be mild, resulting from a smooth transition from local perturbations. On the other hand, when the expressive power of the estimator is excessive, we expect that local perturbations can cause significant shifts in performance, leading to a larger generalization error. Therefore, even though methods like neural ODE [9] could nicely interpolate a given dynamical system, they typically perform poorly in extrapolation. We find that SINDy effectively avoids this issue, as it aims to discover the governing equation in its exact form.
@@ 4 Computational Experiments
In the following, we perform case studies across a range of scientific and engineering problems.
Sea-surface temperature
The first example we consider is that of global sea-surface temperature. The SST data contains 1,400 weekly snapshots of the weekly mean sea surface temperature from 1992 to 2019 reported by NOAA [44]. The data is represented by a 180 $\times$ 360 grid, of which 44,219 grid points correspond to sea-surface locations. We standardize the data with its own min and max, which transforms the sensor measurements to within the numerical range of $(0,1)$. We randomly select 250 sensors from the possible 44,219 locations and set the lag parameter to 52 weeks. The inclusion of 250 sensors is a substantial deviation from previous work with SHRED in which far fewer sensors were used [52]. However, we found greater robustness in the application of E-SINDy to the learned latent state when more sensors were utilized. Thus, for each input-output pair, the input consists of the 52-week trajectories of the selected sensors, while the output is a single temperature field across all 44,219 spatial locations. SINDy-SHRED aims to reconstruct the entire sea surface temperature locations from these randomly selected sparse sensor trajectories. We include the details of the experimental settings of SINDy-SHRED in the Appendix B.1. From the discovered coordinate system, we define the representation of physics - the latent hidden state space - to be $(z_{1},z_{2},z_{3})$. The dynamics progresses forward via the following set of equations:
The analytic solution to this system of ODEs will have the form
The value of $\omega_{1}=1.99\pi$, approximately corresponding to the expected period of 1 year. $\lambda_{1}=0.00763$ indicating a slow decay of the oscillatory mode with half-life $90.84$ years. $\lambda_{3}=0.01527$ indicating global increases in temperature with doubling time $45.39$ years. The explicit solution is given in Appendix E.1.
Figure 3: Extrapolation of latent representation in SINDy-SHRED from the discovered dynamical system for SST. Colored: true latent representation. Grey: SINDy extrapolation.
Figure 4: Decoder reconstruction of three independent directions $[1,0,0]$, $[0,1,0]$ and $[0,0,1]$ in the latent space.
The discovery of a linear system describing the evolution of the latent state is in line with prior work on SST data [13] in which it as assumed that the underlying physics is an advection-diffusion PDE. In Fig. 3 (a) we further present the accuracy of this discovered system by forward simulating the system from an initial condition for a total of 27 years (c.f. Fig. 17). It is observable how the discovered law is close to the true evolution of latent hidden states and, critically, there appears to be minimal phase slipping. Extrapolating the latent state space via forward integration, we can apply the shallow decoder to return forecasts of the high-dimensional data. Doing so, we find an averaged MSE error of $0.57\pm 0.10^{\circ}C$ for all prediction lengths in the test dataset. In Fig. 5, we show SINDy-SHRED produces stable long-term predictions for SST data. We further include Fig. 18 to demonstrate the extrapolation of each sensor. The sensor level prediction is based on the global prediction of the future frame, and we visualize the signal trajectory of specific sensor locations. We find SINDy-SHRED is robust for out-of-distribution sensors, though its extrapolation may not accurately capture anomalous events.
Figure 5: Long-term global sea-surface temperature prediction via SINDy-SHRED from week 0 to week 100. We crop the global temperature map for better visualization.
3D Atmospheric ozone concentration
The atmospheric ozone concentration dataset [2] contains a one-year simulation of the evolution of an ensemble of interacting chemical species through a transport operator using GEOS-Chem. The simulation contains 1,456 temporal samples with a timestep of 6 hours over one year for 99,360 (46 by 72 by 30) spatial locations (latitude, longitude, elevation). The data presented in this work has compressed by performing an SVD and retaining only the first 50 POD modes. As with the SST data, we standardize the data within the range of $(0,1)$ and randomly select and fix 3 sensors out of 99,360 spatial locations ($0.5\%$). We include the details of the experimental settings of SINDy-SHRED in the Appendix B.2. The converged latent representation presents the following SINDy model:
The analytic solution to this system of ODEs will have the form
[…]
where $\omega_{1}=0.0079,$ $\lambda_{1}=0.003$, and $\lambda_{3}=0.003.$ The complete solution specifying $\mathbf{v}_{i}$ and $c_{i}$ is given in Appendix E.2.
Figure 6: Extrapolation of latent representation in SINDy-SHRED from the discovered dynamical system for Ozone data. Colored: true latent representation. Grey: SINDy extrapolation.
Figure 7: Long-term global Ozone data prediction via SINDy-SHRED with elevation 0 from week 0 to week 14.
GoPro physics data: flow over a cylinder
@@
In this subsection, we demonstrate the performance of SINDy-SHRED on an example of so-called “GoPro physics modeling.” The considered data is collected from a dyed water channel to visualize a flow over a cylinder [1]. The Reynolds number is 171 in the experiment. The dataset contains 11 seconds of video taken at 30 frames per second (FPS). We manually perform data augmentation and repeat the latter part of the video once to increase the number of available training samples . We transfer the original RGB channel to gray scale and remove the background by subtracting the mean of all frames. After the prior processing step, the video data has only one channel (gray) within the range $(0,1)$ with a height of 400 pixels and a width of 1,000 pixels. We randomly select and fix 200 pixels as sensor measurements from the entire $400,000$ space, which is equivalent to only $0.05\%$ of the data. We set the lag parameter to 60 frames. We include the details of the experimental settings of SINDy-SHRED in the Appendix B.4.
@@
where $\omega_{1}=1.52$, $\omega_{3}=1.05$, $\lambda_{1}=0.01$, $\lambda_{3}=0.11$, and $\lambda_{5}=-0.20$. The complete explicit solution is given in E.3.
@@
Compared to the systems discovered in all previous examples, the flow over a cylinder model is much more complex with significant nonlinear interactions. In Eqn. 16, we find that $z_{1}$ and $z_{3}$ behave like a governing mode of the turbulence swing; $z_{2}$ and $z_{4}$ further depict more detailed nonlinear effects. We further present a linear model derived from Koopman-SHRED in Eqn. 17 with its latent space evolution demonstrated in Fig. 10. We show the result of extrapolating this learned representation. We generate the trajectory from the initial condition at time point 0 and perform forward integration for extrapolation. As shown in Fig. 8, the learned ODE closely follows the dynamics of $z_{1}$ and $z_{3}$ up to 7 seconds (210 timesteps); $z_{2}$ and $z_{4}$ also have close extrapolation up to 3 seconds. The Koopman-SHRED model closely follows the trend but deviates more significantly from the governing dynamics.
@@
This learned representation nicely predicts the future frames in pixel space. In SINDy-SHRED, the shallow decoder prediction has an averaged MSE error of $0.030$ (equivalently $3\%$) over the entire available trajectory. Koopman-SHRED has relatively worse prediction error with an averaged MSE error of $0.053$. In Fig. 9, we observe that the autoregressively generated prediction frames closely follow the true data, and further in Fig. 23, we find that the predictions are still stable after 1,000 frames, which is out of the size of the original dataset. The sensor-level prediction in Fig. 22 further demonstrates the accuracy of reconstruction in detail.
[…]
Figure 8: (top panel) Extrapolation of latent representation in SINDy-SHRED from the discovered dynamical system for flow over a cylinder data. Colored: true latent representation. Grey: SINDy extrapolation. (bottom panel) Evolution of dynamical system (16). The left two panels are the phase plane $z_{1}$ vs $z_{2}$ and $z_{3}$ vs $z_{4}$ respectively. The magenta plus symbols are the fixed points. The nonlinear limit cycle behavior is shown in the right panel for $z_{j}(t)$.
Figure 9: Long-term pixel space video prediction via SINDy-SHRED. We demonstrate the forward prediction outcome up to 180 frames.
Figure 10: Extrapolation of latent representation in SINDy-SHRED from the discovered dynamical system for flow over a cylinder data. Colored: true latent representation. Grey: SINDy extrapolation.
Baseline study: prediction of single shot real pendulum recording
@@
In this subsection, we compare the performance of SINDy-SHRED to other popular existing learning algorithms. We perform the baseline study particularly on video data of a pendulum since many deep learning algorithms are hard to scale up to deal with large scientific data. In the following, we demonstrate the result of video prediction on the pendulum data using ResNet [23], convolutional LSTM (convLSTM) [48], and PredRNN [51], and SimVP [20]. The pendulum in our experiment is not ideal and includes complex damping effects. We use a nail on the wall and place the rod (with a hole) on the nail. This creates complex friction, which slows the rod more when passing the lowest point due to the increased pressure caused by gravity. The full model we discovered from the video (as shown in Fig. 11) includes four terms:
$\ddot{z}=0.17\dot{z}^{2}-0.06\dot{z}^{3}-10.87\sin(z)+0.48\sin(\dot{z}).$ (19)
@@
As shown in Table 1, SINDy-SHRED outperforms all baseline methods for total error and long-term predictions. Generally, all baseline deep learning methods perform well for short-term forecasting, but the error quickly accumulatesfor longer-term predictions. This is also observable from the prediction in the pixel space as shown in Fig. 12. SINDy-SHRED is the only method that does not produce collapsed longer-term predictions. Interestingly, we observe that incorporating nonlinear terms in the SINDy library is crucial in this case, as Koopman-SHRED struggles to accurately reconstruct the finer details. We attribute this to the fact that strong linear regularization may over-regularize and oversimplify the complex dynamics, leading to a reduction in predictive accuracy. Similar behaviors for Koopman-SHRED are observed when encoding the continuous spectrum using an additional shallow network, as detailed in the Appendix. In Fig. 24, the sensor level prediction also demonstrates the robustness of the SINDy-SHRED prediction. PredRNN is the second best method as measured by the total error. However, PredRNN is expensive in computation which includes a complex forward pass with an increased number of parameters. It is also notable that the prediction of PredRNN collapses after 120 frames, after which only an averaged frame over the entire trajectory is predicted. ConvLSTM has a relatively better result in terms of generation, but the long-term prediction is still inferior compared to SINDy-SHRED. Additionally, we note that 2D convolution is much more computationally expensive. For larger spatiotemporal domains (e.g. the SST example and 3D ozone data), the computational complexity of convolution will scale up very quickly, which makes the algorithm impractical to execute. Similar computational issues will occur for diffusion models and generative models, which is likely to be impractical to compute, and unstable for longer-term predictions. In summary, we observe that SINDy-SHRED is not only a more accurate long-term model, but is also faster to execute and smaller in size.
@@
Figure 11: Extrapolation of latent representation in SINDy-SHRED from the discovered dynamical system for the pendulum moving data. Blue: true latent representation. Grey: SINDy extrapolation.
Figure 12: The pendulum video generation outcome from ResNet, SimVP, ConvLSTM, PredRNN, and SINDy-SHRED from frame 20 to frame 245.

### v2

@@ 3.2.2 Latent space regularization via ensemble SINDy
We first note that we deviate from the original SHRED architecture by using a GRU as opposed to an LSTM. This choice was made because we generally found that GRU provides a smoother latent space and propagates only a single hidden state. Now, recall that our goal is to find a SHRED model with a latent state that is within the SINDy-class functional. However, the initial latent representation from SHRED does not follow the SINDy-based ODE structure at all. On the one hand, if we naively apply SINDy to the initial latent representation, the discovery is unlikely to fit the latent representation trajectory. On the other hand, if we directly replace the GRU unit to SINDy and force the latent space to follow the discovered SINDy model, it might lose information that is important to reconstruction the entire spatial domain. Therefore, it is important to let the two latent spaces align progressively.
In Algorithm 1, we describe our training procedure that allows the two trajectories to progressively align with each other. To further ensure a gradual adaptation and avoid over-regularization, we introduce ensemble SINDy units with varying levels of sparsity constraints, which range in effect from promoting a full model (all terms in the library are active) to a null model (where no dynamics are represented). From the initial latent representation ${\bm{z}}^{\text{iter 0}}_{1:t}$ from SHRED, the SINDy model first provides an initial estimate of ensemble SINDy coefficients $\{\hat{\Xi}_{0}^{i}\}_{i=b}^{B}$. Then, the parameters of SHRED will be updated towards the dynamics simulated by $\{\hat{\Xi}_{0}\}_{i=b}^{B}$, which generates a new latent representation trajectory ${\bm{z}}^{\text{iter 1}}_{1:t}$. We iterate this procedure and jointly optimize the following loss function to let the SHRED latent representation trajectory approximate the SINDy generated trajectory:
@@ 3.2.3 Latent space linearization via the Koopman operator
Koopman operator theory [39, 40] provides an alternative approach to solving these problems by linearizing the underlying dynamics. The linearized embedding is theoretically grounded to be able to represent nonlinear dynamics in a linear framework, which is desirable for many applications in science and engineering, including control, robotics, weather modeling, and so on. From the transformed measurements ${\bm{z}}=g({\bm{x}})$ from the true system ${\bm{x}}$, the Koopman operator $\mathcal{K}$ is an infinite-dimensional linear operator given by
@@
To enable continuous spectrum for better approximation for the Koopman operator, one could adapt an additional neural network to learn the eigenvalues $\lambda_{i}$’s and form the linear dynamics $K$ from the learned eigenvalues [50]. We also provide detailed implementation and experimental results in the Appendix. This flexibility allows the system to learn dynamical systems in a more general setting, accommodating various initial conditions and experimental setups. In our experimental study, we observe that this strategy may perform comparably to learning a fixed ${\bm{K}}$ for linear dynamics under fixed experimental environments.
3.3 Theoretical analysis on dynamical system learning
Suppose that the dynamical system has the form $\dot{{\bm{x}}}=f({\bm{x}})$, and we have measurements of
${\bm{x}}=\{{\bm{x}}_{1},{\bm{x}}_{2},...,{\bm{x}}_{t},{\bm{x}}_{t+1},...,{\bm{x}}_{T}\},$ (8)
@@ 3.3 Theoretical analysis on dynamical system learning
with time gap $\Delta t$. The empirical loss function $L_{n}(\cdot)$ of a function $f$ given the empirical distribution $P_{n}$ which consists of data samples $\mathcal{D}=\{({\bm{x}}_{i},\dot{{\bm{x}}_{i}})\}_{i=1}^{n}$ is
$L_{n}\left(f\right)=\frac{1}{n}\sum_{i=1}^{n}\ell\left(f({\bm{x}}_{i}),{\bm{x}}_{i}\right),$ (9)
while the true loss is $L\left(f\right)=\mathbb{E}\left[\ell\left(f(x),y\right)\right]$.
@@
From empirical process theory, there is a generalization gap between $L_{n}(f)$ and $L(f)$, which is correlated with the expressive power of the functional class [74, 72]. When the expressive power of the estimator functional class is small, the generalization error is expected to be mild, resulting from a smooth adaptation from local perturbations. On the other hand, when the expressive power of the estimator is excessive, we expect that local perturbations can cause significant shifts in performance, leading to a larger generalization error.
Therefore, even though neural networks are universal approximators [36], they can perform poorly in extrapolation. SINDy effectively avoids this issue because the functional class is moderately sized, and the continuous formulation allows SINDy to reach a much lower error. We present the following theorems to compare error bounds for dynamical system learning between the SINDy-class and the neural network class. First, for a SINDy-class, we have the following theorem.
Theorem 1.
Suppose we have a SINDy-class functional with a library of functions $\Theta(x)$, and we estimate the coefficient vector $\xi$ by $\hat{\xi}$ using least squares. Let $\xi^{*}$ be the true coefficient vector. Under regularity conditions in the Appendix, we have the expected error in predicting the dynamical system after time $T$ is bounded by
$E||\hat{x}(T)-x(T)||\leq\mathcal{O}\left(e^{LT}s\sqrt{\frac{p}{n}}\right)$ (10)
where $L$ is the Lipschitz constant of the system, $s$ is the level of noise, $p$ is the number of functions in the library, and $n$ is the number of samples.
Furthermore, with probability $1-\delta$, the error in predicting the dynamical system after time $T$ is bounded by
$||\hat{x}(T)-x(T)||\leq\mathcal{O}\left(e^{LT}s\sqrt{\frac{p}{n}\log\left(\frac{1}{\delta}\right)}\right).$ (11)
Similarly, for the neural network class of functions, we have:
Theorem 2.
Suppose we have a neural network functional with $k$ layers of ReLU activation functions and parameters $\theta=\left({\bm{W}}_{1},\dots,{\bm{W}}_{k}\right)$, which computes functions
$f({\bm{x}};\theta)=\sigma_{k}\left({\bm{W}}_{k}\sigma_{k-1}\left({\bm{W}}_{k-1}\cdots\sigma_{1}\left({\bm{W}}_{1}{\bm{x}}\right)\right)\right).$ (12)
Assume that the weights are bounded such that $\left\|{\bm{W}}_{i}\right\|_{F}\leq B$ for all $i$. Then, with probability at least $1-\delta$, the error in predicting the dynamical system after $H=T/\Delta t$ steps is bounded by
$||\hat{x}(T)-x(T)||\leq\mathcal{O}\left((\log n)^{4}B^{k(H+1)}\sqrt{\frac{k}{n}}+\frac{\log(1/\delta)}{n}\right),$ (13)
where $n$ is the number of samples.
From the error bounds above, we notice that compared to traditional learning problems (e.g. classification, regression), dynamical system learning is more challenging as the error grows exponentially with increasing time horizon. Short-term local error accumulates as the differential equation steps forward, leading to a large error for longer-time extrapolation. This exponential error behavior exists even for linear problems. From the error bound, it becomes evident why neural networks may encounter significant challenges in learning such problems. The challenge is mainly due to the presence of the term $B^{kH}$. The bound on weights $B$ is expected to increase significantly for complex dynamical systems that need to model large local temporal derivatives. And, the number of layers $k$ is expected to increase for more complex and chaotic dynamics. To generate high-quality video sequences, finer temporal steps are required for a higher frame-per-second rate, which further exacerbates error growth. Therefore, the $B^{kH}$ term could be extremely large in practice.
To control these errors contributed by $B^{kH}$, one needs to collect exponentially more samples to ensure the generalization error is small. This explains the observation that many video generation methods relying purely on neural networks struggle to follow physical laws accurately, as the required sample size is prohibitively large. Transformer networks tend to memorize example trajectories in their key-value pairs, but its mechanism is not inherently designed to capture the exact physical law, which can often produce poor extrapolations from unseen initial conditions [85].
@@ 4 Computational Experiments
In the following, we perform case studies across a range of scientific and engineering problems. We begin with building models directly from video data, which includes flow around a cylinder and a pendulum. We then consider real-world sea-surface temperature data and conclude with turbulent flow examples.
4.1 GoPro physics video data: flow over a cylinder
@@
In this subsection, we demonstrate the performance of SINDy-SHRED on an example of so-called “GoPro physics modeling.” The considered data is collected from a dyed water channel to visualize a flow over a cylinder [1]. The Reynolds number is 171 in the experiment. The dataset contains 11 seconds of video taken at 30 frames per second (FPS). We transfer the original RGB channel to gray scale and remove the background by subtracting the mean of all frames. After the prior processing step, the video data has only one channel (gray) within the range $(0,1)$ with a height of 400 pixels and a width of 1,000 pixels. We randomly select and fix 200 pixels as sensor measurements from the entire $400,000$ space, which is equivalent to only $0.05\%$ of the data. We set the lag parameter to 60 frames. We include the details of the experimental settings of SINDy-SHRED in the Appendix D.3.
@@
[…]
Figure 2: (a) Extrapolation of latent representation in SINDy-SHRED from the discovered dynamical system for flow over a cylinder data. Colored: true latent representation. Grey: SINDy extrapolation. (b) Evolution of dynamical system (14). The left two panels are the phase plane $z_{1}$ vs $z_{2}$ and $z_{3}$ vs $z_{4}$ respectively. The magenta plus symbols are the fixed points. The nonlinear limit cycle behavior is shown in the right panel for $z_{j}(t)$. (c) Long-term pixel space video prediction via SINDy-SHRED. We demonstrate the forward prediction outcome up to 180 frames.
@@
where $\omega_{1}=1.52$, $\omega_{3}=1.05$, $\lambda_{1}=0.01$, $\lambda_{3}=0.11$, and $\lambda_{5}=-0.20$. The complete explicit solution is given in G.3.
@@
Compared to the systems discovered in all previous examples, the flow over a cylinder model is much more complex with significant nonlinear interactions. In Eqn. 14, we find that $z_{1}$ and $z_{3}$ behave like a governing mode of the turbulence swing; $z_{2}$ and $z_{4}$ further depict more detailed nonlinear effects. We further present a linear model derived from Koopman-SHRED in Eqn. 15 with its latent space evolution demonstrated in Fig. 19. We show the result of extrapolating this learned representation. We generate the trajectory from the initial condition at time point 0 and perform forward integration for extrapolation. As shown in Fig. 2 (a), the learned ODE closely follows the dynamics of $z_{1}$ and $z_{3}$ up to 7 seconds (210 timesteps); $z_{2}$ and $z_{4}$ also have close extrapolation up to 4 seconds. The Koopman-SHRED model closely follows the trend (in Fig. 19) but deviates more significantly from the governing dynamics.
@@
This learned representation nicely predicts the future frames in pixel space. In SINDy-SHRED, the shallow decoder prediction has an averaged MSE error of $0.030$ (equivalently $3\%$) over the entire available trajectory. Koopman-SHRED has relatively worse prediction error with an averaged MSE error of $0.053$. In Fig. 2 (c), we observe that the autoregressively generated prediction frames closely follow the true data, and further in Fig. 21, we find that the predictions are still stable after 1,000 frames, which is out of the size of the original dataset. The sensor-level prediction in Fig. 20 further demonstrates the accuracy of reconstruction in detail.
4.2 Prediction and baseline study of single shot pendulum video
@@
In this subsection, we compare the performance of SINDy-SHRED to existing state-of-the-art learning algorithms. In the following, we demonstrate the result of video prediction on the pendulum data using ResNet [33], convolutional LSTM (convLSTM) [67], and PredRNN [76], and SimVP [29]. The pendulum in our experiment is not ideal and includes complex damping effects. We use a nail on the wall and place the rod (with a hole) on the nail. This creates complex friction, which slows the rod more when passing the lowest point due to the increased pressure caused by gravity. The full model we discovered from the video (as shown in Fig. 3 (a)) includes four terms:
$\ddot{z}=0.17\dot{z}^{2}-0.06\dot{z}^{3}-10.87\sin(z)+0.48\sin(\dot{z}),$ (17)
which includes complex damping effects via $\sin(\dot{z}),\dot{z}^{2},\dot{z}^{3}$, in contrast to the commonly assumed linear damping.
@@
As shown in Table 1, SINDy-SHRED outperforms all baseline methods for total error and long-term predictions. Generally, all baseline deep learning methods perform well for short-term forecasting, but the error quickly accumulates for longer-term predictions. This is also observable from the prediction in the pixel space as shown in Fig. 3 (b). SINDy-SHRED is the only method that does not produce collapsed longer-term predictions. Interestingly, we observe that incorporating nonlinear terms in the SINDy library is crucial in this case, as Koopman-SHRED struggles to accurately reconstruct the finer details. We attribute this to the fact that strong linear regularization may over-regularize and oversimplify the complex dynamics, leading to a reduction in predictive accuracy. Similar behaviors for Koopman-SHRED are observed when encoding the continuous spectrum using an additional shallow network. In Fig. 24, the sensor level prediction also demonstrates the robustness of the SINDy-SHRED prediction. PredRNN is the second best method as measured by the total error. However, PredRNN is expensive in computation which includes a complex forward pass with an increased number of parameters. It is also notable that the prediction of PredRNN collapses after 120 frames, after which only an averaged frame over the entire trajectory is predicted. ConvLSTM has a relatively better result in terms of generation, but the long-term prediction is still inferior compared to SINDy-SHRED. Additionally, is should be noted that 2D convolution is much more computationally expensive. For larger spatiotemporal domains (e.g. the SST example and 3D ozone data), the computational complexity of convolution will scale up very quickly, which makes the algorithm nearly impossible to deploy in practice. Similar computational issues will occur for diffusion models and generative models, which is likely to be impractical to compute, and unstable for longer-term predictions. In summary, we observe that SINDy-SHRED is not only a more accurate long-term model, but is also faster to execute and smaller in size.
@@
Figure 3: (a) Extrapolation of latent representation in SINDy-SHRED from the discovered dynamical system for the pendulum moving data. Blue: true latent representation. Grey: SINDy extrapolation. (b) The pendulum video generation outcome from ResNet, SimVP, ConvLSTM, PredRNN, and SINDy-SHRED from frame 20 to frame 245.